Determine the quantity of heat required to convert 1kg of ice formally at -20°c to water at 100°c and steam at 100°c
1
Expert's answer
2020-01-28T13:51:07-0500
"Q_w=m(c_i(T_2-T_1)+c_w(T_3-T_2)+L_f)"

"Q_w=1(2.108(0-(-20))+4.187(100-0)+334)=795\ kJ"

"Q_s=Q_w+mL_v=795+1(2265)=3060\ kJ"